OpenStudy (meilendurcer):

Which of the following shows the real number expressions in order from least to greatest? answers below

OpenStudy (meilendurcer):

\[6-\sqrt{8}\],\[3^{2}-2\],\[\sqrt{16}+0.8\]

OpenStudy (agent0smith):

You'll be better off evaluating them before comparing them, I think. \[6-\sqrt{8}= \] \[3^{2}-2 = \] \[\sqrt{16}+0.8 = \] Compare the results.